# Streamlining Redundant Layers to Compress Large Language Models

### v3

@@ 4.3 Baseline
4.4 Main Results
We present accuracy and stability for different methods on the classification benchmarks in Table 2 and Table 3, respectively, and Table 4 for the generative benchmarks. The results demonstrate that our proposed LLM-Streamline consistently outperforms the baseline methods, effectively retaining the majority of the capabilities of LLMs across various tasks. Specifically, in classification tasks, on Llama2-7B, LLM-Streamline surpasses SliceGPT by 7% in terms of accuracy and 15% in terms of stability. LLM-Streamline also surpasses LaCo by 5% in terms of accuracy. For generation tasks, LLM-Streamline retains nearly 77% of Llama2-7B’s capabilities and 94% of OPT-6.7B’s capabilities, significantly outperforming other pruning methods. We find that almost all of the pruning methods fail on the GSM8K dataset, which reflects the difficulty of math word problems.
Additionally, comparing the average stability (Ave.) in Table 3 with the retrained performance (Pre.) in Table 2 across different pruning methods reveals that stability is often much lower than accuracy. We also observe that accuracy on Race-M and Race-H even increases after model pruning. This suggests that pruned models tend to make correct guesses on some classification questions that they are uncertain, highlighting the limitations of accuracy as a sole measure of pruning method performance. We also conduct experiments on smaller models, such as OPT-1.3B and OPT-2.7B. Details can be found in Appendix  C.1.
Accuracy 45.1 45.2 45.6 44.4 46.0 45.8 43.8 44.2
Stability 81.2 82.7 81.9 79.2 80.7 83.7 82.6 83.3
[…]
Figure 4: (a) Stability of the pruned Llama-7B at pruning ratios of about 2%, 11%, 20%, and 25%. Since stability is calculated solely for compressed model compared to the original model, we could not compute the stability for TinyLlama-1.1B and OpenLlama-3B. (b) Accuracy of the pruned LLama-7B at the same pruning ratios. Metrics are averaged across the classification benchmarks.
4.5 Impact of Different Lightweight Networks
To investigate the impact of the model size, architecture, and initialization methods of lightweight networks on the pruned model’s performance, we conduct experiments on Llama2-7B using the different architecture as lightweight network, including Feed-Forward Neural Network (FFN), SwiGLU-based Feed-Forward Neural Network (SwiGLU), and Transformer Layer (Layer). We also explore various initialization methods for the Transformer Layer, including random initialization (Layer-Random), inheritance of the first pruned layer (Layer-First), inheritance of the last pruned layer (Layer-Last), and averaging the pruned layers (Layer-Avg). The average accuracy and stability metrics across all the classification benchmarks are presented in Table 5, with detailed results on each benchmark available in Appendix C.2. The results show that FFN achieves the best results. Meanwhile, for the Transformer Layer, inheritance of the pruned first layer yields the best results. In addition, We plot the validation loss curves during the training process for different lightweight networks, as shown in Fig. 3. We can observe that FFN and SwiGLU have already converged by the 10th epoch, whereas the loss of Transformer Layer is still decreasing. This indicates that the Transformer layer still has potential, and further training could yield better results, but this would require more computing resources.
4.6 Impact of Different Pruning Ratios
To verify the performance of the model at other pruning ratios, in addition to the about 25% pruning ratio, we also evaluate our method’s performance at other pruning ratios on Llama2-7B, with pruning ratios 2%, 11%, and 20%. The average stability and accuracy metrics across all the classification benchmarks are shown in Fig. 4, with details on each benchmark presented in Appendix C.3. By comparing the performance of Llama2-7B, TinyLlama-1.1B, OpenLlama-3B-v2, and the model at different pruning ratios, we approximate that the performance of the pruned model is linearly correlated with the number of parameters, demonstrating the effectiveness of our method.
4.7 Comparison of Layer Replacement and LoRA
We compare the performance of layer replacement with LoRA. Since layer replacement is trained based on hidden states with a different training objective than LoRA, we additionally train one epoch using the language model loss for layer replacement when comparing it with LoRA. The training details can be found in the Appendix B.2. For layer replacement, we freeze the original model’s weights and only train the lightweight network. For LoRA, we set the rank to 384 to ensure that the number of parameters involved in training is approximately the same as that of lightweight networks. We randomly extract 30,000 samples from SlimPajama-6B as training data. The average accuracy and stability across all the classification benchmarks are presented in Table 6, with detailed results for each benchmark available in Appendix C.4. The results show that layer replacement outperforms LoRA in both accuracy and stability. In addition to the impact on performance, the post-training process for layer replacement also demands significantly less GPU memory than LoRA requires.
Layer-First Layer-Last Layer-Avg FFN FFN SwiGLU SwiGLU LoRA
Accuracy 46.7 46.8 46.2 45.8 46.3 44.4 45.5 44.6
Stability 85.7 85.6 83.9 83.4 85.2 84.7 84.7 82.6
GPU Memory (G) 27.8 27.8 27.8 25.6 27.0 25.3 26.4 47.6
[…]
5 Related Work
Model pruning can be categorized into unstructured pruning and structured pruning. Unstructured pruning compresses the model by converting its dense matrices into sparse matrices, retaining the dense matrix, and is more hardware-friendly louizos2017l0 ; frantar2023sparsegpt ; das2023beyond ; sun2023simple . Structured pruning maintains performance at high compression rates but requires specialized sparse operators, limiting its applicability chen2023lorashear ; xia2023sheared ; ashkboos2024slicegpt ; ma2023llmpruner .
LLM-Prunerma2023llmpruner is a structured pruning method that uses gradients and activations to estimate the importance of groups of modules in LLMs and prunes the less important ones. SliceGPTashkboos2024slicegpt , another structured pruning method, inserts dimensionality reduction matrices into LLMs and uses PCA to initialize and eliminate unimportant components within these matrices. LaCoyang2024laco prunes layers by merging similar layers. Unlike LaCo, our proposed method, LLM-Streamline, uses a lightweight network to compensate for the loss caused by the pruned layers instead of merging them.
6 Conclusion
This work introduces a layer pruning-and-replacement algorithm for LLMs. We identify shortcomings in the existing accuracy metric and propose a new metric called “stability” for model compression. Extensive experiments show that this layer replacement method using a lightweight network outperforms previous state-of-the-art pruning methods.
Appendix A Data Distribution
We extract the training data from different domains based on the data distribution strategy proposed in Sheared-LLaMa [17]. The detailed data distribution is shown in Table 7.
CC GitHub Book StackExchange Wiki ArXiv C4
SlimPajama-6B 54.1% 4.2% 3.7% 2.8% 3.1% 3.4% 28.7%
Ours 36.1% 0.8% 9.1% 1.0% 3.1% 0.7% 49.2%
Appendix B Training Implementation Details
B.1 Lightweight Network Training Details
B.2 Post Training Details
For layer replacement, in order to have a fairer comparison with LoRA, we conduct one epoch of post-training with a learning rate of 5e-5, a weight decay of 1e-3, and a batch size of 32. This process takes less than an hour on a single A800 GPU. For LoRA, the model is trained over 10 epochs with a learning rate of 5e-5, a weight decay of 1e-3, and a batch size of 32, requiring approximate 10 hours on a single A800 GPU.
Appendix C Detailed Experimental Results
C.1 Experimental Results of OPT-1.3B and OPT-2.7B
@@ Appendix C Detailed Experimental Results
We also conduct experiments on small models (OPT-1.3B and OPT-2.7B). The experimental results are shown in Table 8. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method, SliceGPT, across different pruning rates.
LLM Method Ratio Benchmarks Ave. Per.
@@
C.2 Detailed Results of Different Lightweight Networks
@@
LLM Method Ratio Benchmarks Ave. Per.
@@
LLM Method Ratio Benchmarks Ave.
@@
C.3 Detailed Results of Different Pruning Ratio
The detailed results of accuracy and stability on LLMs under different pruning ratios on differnet classification benchmarks are shown in Table 11 and Table 12. The experiment results show that the performance of the pruned model is linearly correlated with the number of parameters, demonstrating the effectiveness of our method.
LLM Method Ratio Benchmarks Ave. Per.
@@ C.3 Detailed Results of Different Pruning Ratio
LLM Method Ratio Benchmarks Ave.
@@
C.4 Detailed Results of Layer Replacement and LoRA
@@
LLM Method Ratio Benchmarks Ave. Per.
@@
LLM Method Ratio Benchmarks Ave.

### v4

@@ Appendix D Detailed Experimental Results
Table 13: Accuracy of different pruning methods on generation benchmarks by pruning OPT-6.7B.
D.2 Experimental Results of Baichuan-7B and Baichuan-13B
@@
We also conduct experiments on Baichuan-7B and Baichuan-13B. The experimental results are shown in Table 14, Table 15 and Table 16. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
LLM Method Ratio Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Baichuan-7B Dense 0.00% 55.8 35.3 91.3 61.4 39.4 58.4 65.3 77.6 29.5 30.4 43.7 43.8 52.7 100
LLMPruner 24.2% 43.7 33.9 65.9 40.5 36.5 48.2 52.2 68.1 22.6 22.0 24.2 25.3 40.2 76.3
Ours (None) 24.2% 33.2 32.7 25.8 60.8 36.5 36.0 34.6 58.7 22.1 21.5 25.7 38.8 35.5 67.4
Ours (FFN) 25.1% 53.1 36.3 69.4 53.1 36.5 48.7 53.2 69.4 23.2 24.5 37.7 39.1 45.4 86.1
Ours (Layer) 24.2% 55.0 36.0 77.4 48.1 36.5 49.8 54.3 69.0 22.9 23.8 39.8 41.1 46.1 87.5
Baichuan-13B Dense 0.00% 61.5 36.4 91.5 65.8 49.0 64.2 69.1 78.2 48.1 46.0 54.8 55.3 60.0 100
Ours (None) 24.7% 48.8 34.8 50.2 62.2 40.4 46.4 56.7 68.2 30.6 27.7 52.9 55.1 47.8 79.7
Ours (FFN) 25.5% 58.3 35.1 77.5 64.1 36.5 57.7 58.2 69.4 26.5 28.8 53.1 54.3 51.6 86.0
Ours (Layer) 24.7% 59.1 36.1 83.7 62.0 36.5 58.2 59.4 71.8 27.8 25.0 52.3 56.1 52.3 87.2
Table 14: Accuracy of different pruning methods on classification benchmarks by pruning Baichuan-7B and Baichuan-13B.
LLM Method Ratio Benchmarks Average
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Baichuan-7B LLMPruner 24.2% 70.2 40.0 71.4 24.9 91.4 75.3 75.9 82.4 68.8 66.8 54.7 53.6 64.6
Ours (None) 24.2% 55.5 45.1 31.3 76.8 93.3 64.7 50.1 61.0 68.9 65.9 74.0 60.8 62.3
Ours (FFN) 25.1% 84.1 77.8 75.6 65.4 93.3 76.5 83.1 84.3 77.1 70.3 74.2 74.6 78.0
Ours (Layer) 24.2% 86.3 79.3 82.3 40.7 93.3 77.6 81.3 83.5 75.8 71.0 75.9 75.6 76.9
Baichuan-13B Ours (None) 24.7% 67.2 75.8 54.1 66.6 51.0 74.9 74.6 78.5 52.3 55.2 82.7 89.3 68.5
Ours (FFN) 25.5% 85.7 87.3 82.1 81.3 43.3 81.2 81.7 79.8 46.6 61.1 84.3 83.7 74.8
Ours (Layer) 24.7% 88.6 92.7 89.3 72.9 43.3 83.5 86.1 88.2 49.2 52.1 83.4 90.4 76.6
Table 15: Stability of different pruning methods on classification benchmarks by pruning Baichuan-7B and Baichuan-13B.
LLM Method Ratio Benchmarks Average RP
Xsum GSM8K StrategyQA
Baichuan-7B Dense 0.00% 19.1 9.84 55.5 28.1 100
LLMPruner 24.2% 12.6 1.74 40.4 18.2 64.8
Ours (None) 24.2% 0.3 0 0 0.1 0
Ours (FFN) 25.1% 19.3 2.11 41.1 20.8 74.0
Ours (Layer) 24.2% 18.2 1.36 38.7 19.4 69.0
Baichuan-13B Dense 0.00% 24.6 27.1 61.1 37.6 100
Ours (None) 24.7% 2.1 1.2 12.3 5.2 13.8
Ours (FFN) 25.5% 23.1 2.1 47.3 24.2 64.4
Ours (Layer) 24.7% 22.2 2.4 43.2 22.6 60.1
Table 16: Accuracy of different pruning methods on generation benchmarks by pruning Baichuan-7B and Baichuan-13B.
D.3 Experimental Results of Baichuan2-7B and Baichuan2-13B
We also conduct experiments on Baichuan2-7B and Baichuan2-13B. The experimental results are shown in Table 17, Table 18 and Table 19. The results indicate that our proposed LLM-Streamline is superior to the concurrent SOTA method, LaCo.
LLM Method Ratio Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Baichuan2-7B Dense 0.00% 64.4 33.4 85.5 63.1 42.3 63.1 67.6 76.1 51.1 52.5 54.7 57.1 59.2 100
LLMPruner 24.2% 39.9 33.9 70.6 50.0 42.3 38.7 52.7 70.4 22.3 22.8 24.9 24.9 41.1 69.4
LaCo* 24.2% 50.9 33.0 76.2 56.2 42.3 47.3 52.3 68.5 27.7 29.0 31.5 31.2 45.5 76.9
Ours (None) 24.2% 45.7 33.0 58.0 62.6 36.5 41.9 46.3 62.4 25.6 27.4 43.0 46.5 44.1 74.5
Ours (FFN) 25.1% 58.2 33.0 74.1 61.2 36.5 47.6 54.3 68.0 29.1 30.5 52.1 56.7 50.1 84.6
Ours (Layer) 24.2% 60.4 34.9 72.2 62.7 36.5 48.8 52.5 67.0 35.5 36.8 54.0 56.3 51.5 87.0
Baichuan2-13B Dense 0.00% 65.6 33.2 86.7 66.8 42.3 65.6 71.1 78.1 68.9 67.2 59.6 61.3 63.9 100
LaCo* 24.7% 61.1 33.0 76.7 62.4 44.2 55.5 60.7 68.9 57.8 56.9 51.4 53.7 56.9 89.0
Ours (None) 24.7% 59.1 34.4 81.9 61.8 36.5 53.9 61.9 71.0 63.0 60.4 50.3 57.9 57.7 90.3
Ours (FFN) 25.5% 63.0 33.0 81.7 60.1 36.5 54.7 62.1 70.5 71.1 68.2 57.1 58.2 59.7 93.4
Ours (Layer) 24.7% 63.5 33.0 84.1 62.0 38.5 56.9 63.0 72.0 70.2 66.3 59.1 60.2 60.7 95.0
Table 17: Accuracy of different pruning methods on classification benchmarks by pruning Baichuan2-7B and Baichuan2-13B. “*” indicates that we refer to the results in the original paper.
LLM Method Ratio Benchmarks Average
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Baichuan2-7B LLMPruner 24.2% 62.2 50.6 75.1 55.5 63.5 66.0 80.3 87.0 54.2 51.7 50.0 46.9 61.9
Ours (None) 24.2% 68.5 98.1 63.8 69.1 84.6 61.3 67.3 72.0 50.6 48.0 69.3 74.5 68.9
Ours (FFN) 25.1% 81.1 98.1 77.1 68.2 84.6 66.5 78.1 77.3 61.2 57.7 87.3 89.2 77.2
Ours (Layer) 24.2% 83.3 92.8 74.9 67.6 84.6 68.0 77.2 80.9 64.1 62.9 89.2 88.6 77.8
Baichuan2-13B Ours (None) 24.7% 85.0 88.7 86.5 85.1 82.7 79.5 82.2 85.1 84.6 83.0 74.2 85.6 83.5
Ours (FFN) 25.5% 86.4 99.0 87.2 84.7 82.7 77.6 83.2 85.7 83.2 81.1 90.2 91.7 86.1
Ours (Layer) 24.7% 87.9 99.0 89.0 87.1 84.7 80.2 84.9 86.9 89.4 87.0 91.7 92.5 88.4
Table 18: Stability of different pruning methods on classification benchmarks by pruning Baichuan2-7B and Baichuan2-13B.
LLM Method Ratio Benchmarks Average RP
Xsum GSM8K StrategyQA
Baichuan2-7B Dense 0.00% 21.0 24.8 60.0 35.3 100
LLMPruner 24.2% 14.5 1.4 10.8 8.9 25.2
LaCo* 24.2% 12.0 - - - -
Ours (None) 24.2% 12.1 1.7 30.7 14.8 41.9
Ours (FFN) 25.1% 15.9 2.7 37.1 18.6 52.7
Ours (Layer) 24.2% 16.8 2.3 34.8 18.0 51.0
Baichuan2-13B Dense 0.00% 25.3 53.2 65.9 48.1 100
LaCo* 24.7% 12.3 - - - -
Ours (None) 24.7% 17.2 3.3 37.2 19.2 39.9
Ours (FFN) 25.5% 21.3 3.1 48.8 24.4 50.7
Ours (Layer) 24.7% 20.9 5.5 51.3 25.9 53.8
Table 19: Accuracy of different pruning methods on generation benchmarks by pruning Baichuan2-7B and Baichuan2-13B. “*” indicates that we refer to the results in the original paper.
D.4 Experimental Results of OPT-1.3B and OPT-2.7B
We also conduct experiments on small models (OPT-1.3B and OPT-2.7B). The experimental results are shown in Table 20. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method, across different pruning rates.
LLM Method Ratio Benchmarks Average RP
@@ D.3 Experimental Results of Baichuan2-7B and Baichuan2-13B
D.5 Results of Sufficient Post-training
Following the method outlined in Section 4.7, We conduct experiments using the entire SlimPajama-6B for post-training, and the results are presented in Table 21. As shown, using the entire dataset resulted in a slight improvement, but at a significant computational cost, requiring 100 times the computational time.
LLM Method Training data size Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU Xsum GSM8k StrategyQA
Llama2-7B Dense - 43.8 33.0 41.6 70.8 37.5 66.7 71.3 78.1 33.1 35.5 46.8 31.8 19.4 16.5 60.2 45.7 100.0
Layer-First 30k 43.9 33.0 29.8 70.8 36.5 59.6 64.3 73.4 36.6 37.4 44.9 30.0 19.7 2.05 54.8 42.5 93.0
Layer-First 5.49M 43.5 33.0 33.2 68.8 46.2 61.1 66.5 76.0 31.8 29.9 47.3 31.8 18.2 10.6 58.6 43.8 95.8
Table 21: Detailed accuracy results with different training data volumes.
D.6 Detailed Results of Different Lightweight Networks
@@
LLM Method Ratio Benchmarks Average RP
@@
LLM Method Ratio Benchmarks Average
@@
D.7 Detailed Results of Different Pruning Ratio
The detailed results of accuracy and stability on LLMs under different pruning ratios on different classification benchmarks are shown in Table 24 and Table 25. The experiment results show that the performance of the pruned model is linearly correlated with the number of parameters, demonstrating the effectiveness of our method.
LLM Method Ratio Benchmarks Average RP
@@ D.7 Detailed Results of Different Pruning Ratio
LLM Method Ratio Benchmarks Average
@@
D.8 Detailed Results of Layer Replacement and LoRA
@@
LLM Method Ratio Benchmarks Average RP
@@
LLM Method Ratio Benchmarks Average
@@
Appendix E Limitation
Our method achieves SOTA results compared to existing pruning methods, but its performance still falls short of other commonly used model compression methods, e.g., quantization. Therefore, we plan to enhance the performance of our pruning method and explore combining it with other compression and inference acceleration techniques to make it more practical.